011406:

a post 5m high casts a shadow of 6m.if the shadow of a nearby tree was measured to be 24m,how high is the tree?

Shadow:

The concept you're applying here is that these two shadows are similar, so you can compare them like this: \[\frac{ 5 }{ 6 } = \frac{ x }{ 24 }\]

Shadow:

tl;dr basically just solve for x. The numerators are the height of the object. The denominator is the height of the shade. By comparing them you can use the relationship the height of the shadow has to the object to determine the height of the other object in relation to the height of its shadow.

Shadow:

\[5 \times 24 =6x \rightarrow x = (5 \times 24)/6\]